OpenStudy (jotopia34):

Can anyone assist with the remaining steps in this trig substitution integral: Integral of 2/x^3(sqrtx^2-1) dx Ill post where I'm up to. thanx

OpenStudy (jotopia34):

\[\int\limits_{}\frac{ 2dx }{ x^3\sqrt{x^2-1} }\] I've solved to this point: =∫1+cos(2Ɵ)dƟ

OpenStudy (jotopia34):

\[=∫1+\cos(2Ɵ)dƟ\] =∫1dƟ+∫cos(2Ɵ)dƟ

OpenStudy (anonymous):

I am not sure if I would obtain the same result, which substitution did you choose?. I would say: \[\Large \tan^2\alpha+1=\sec^2\alpha \] so this holds true: \[\Large \sec^2\alpha-1=\tan^2\alpha \] which looks like a perfect substitution for our integral problem: let: \[ \Large x=\sec\alpha\] such that: \[\Large dx=\sec\alpha\tan\alpha d\alpha \]

OpenStudy (anonymous):

I didn't mean to disagree by the way, just had to carry out the steps myself first.

OpenStudy (anonymous):

So it seems like I endup with :\[ \Large 2 \int \cos^2\alpha d\alpha \] if I didn't make any careless mistakes

OpenStudy (anonymous):

and your integration will again lead to: \[\Large \alpha +\sin\alpha\cos\alpha +C\]

OpenStudy (jotopia34):

Not at all, I did the above and I got to that point. then i thought i had to use the sub 1+cos2Ɵ

OpenStudy (anonymous):

that's correct, the two cancels out at this step, so it seems, and you should obtain the same result as I have above, or maybe you have obtained this (double angular expression): \[\Large \alpha + \frac{1}{2}\sin(2\alpha)+C\]

OpenStudy (jotopia34):

The directly above is what I got. Again, I forget how to do the triangle part, to back sub for Ɵ :(

OpenStudy (anonymous):

alright, again we use the relation I posted above: \[\Large \alpha +\sin\alpha\cos\alpha+C\] Because in here it will be easier to back substitute, in the other expression, our result will turn out unpractical and optically displeasing.

OpenStudy (anonymous):

We have performed that substitution: \[\Large x=\sec\alpha=\frac{1}{\cos\alpha} \longrightarrow \cos\alpha=\frac{1}{x}=\frac{ank}{hyp}\]

OpenStudy (jotopia34):

im not following you how 1/cosƟ goes to just cosƟ...?

OpenStudy (anonymous):